OpenStudy (anonymous):

Part 1: Solve the compound inequality 7x - 4 > 31 and 7x - 4 <_ 59 and show all of your work. You can use <= to represent the less than or equal to symbol. Part 2: Use complete sentences to describe the graph of the solution from Part 1

OpenStudy (texaschic101):

7x - 4 > 31 7x - 4 + 4 > 31 + 4 7x > 31 + 4 7x > 35 7x/7 > 35/7 x > 5 7x - 4 <= 59 7x - 4 + 4 <= 59 + 4 7x <= 63 7x/7 <= 63/7 x <= 9 Number line graph.... x is greater then 5 and less then or equal to 9 open circle on 5 shaded to the right, up till nine, with a closed circle on 9

OpenStudy (anonymous):

thank you i think i understand it now